express the perimeter p of a rectangle as function of it sides a and b and compute the perimeter when a=3 and b=5
Answer by bucky(2189)

Two of the characteristics of a rectangle are that it has 4 sides and that its opposite
sides are equal in length. Therefore, if you know that two sides are "a" and "b", then you
also know that there are two other sides of lengths "a" and "b". And you should
know that
the perimeter of a rectangle is found by adding the lengths of all 4 of its sides.
.
So the equation for the perimeter (P) of the given rectangle can be written as:
.
P = a + b + a + b
.
And this equation can be simplified by adding together the common terms on the right side
of the equation to get:
.
P = 2a + 2b
.
This equation gives P, the perimeter, as a function of the sides "a" and "b" of the
rectangle.
.
The problem tells you that a = 3 and b = 5. You can find the perimeter of the rectangle
by substituting these values for "a" and "b" into the equation to get:
.
P = (2*3) + (2*5) = 6 + 10 = 16
.
The answer is that the perimeter of the given rectangle is 16.
.
Hope this helps you to understand the problem.